OSN, then 6. Changing mass is 100 g and mass of the trolley is 500 g friction, acceleration of either mass is how much?
Answer #1

6) T = m (g-a) 7 = 8+ mza milg-a) = f tmea (400) 9-a) = 0.05 + 5.000) a 0.169.8-a) = 0.05 + losa) 0.98-0.1a = 0.05+ o.sa 0.93

Please rate the work if you are satisfied with the solution or else kindly ask your doubts or queries in the comment section below